OpenStudy (anonymous):

ALGEBRA 2: Simplify. I did 6 of these already, this one is confusing me so much! [(t^2 + 6t) ÷ (t - 1)] - [7 ÷ (t - 1)]

OpenStudy (anonymous):

\[(t ^{2}+6t)/(t-1) - 7/(t-1) = (t ^{2}+6t-7)/(t-1)= (t-1)(t+7)/(t-1) = t+7\]

OpenStudy (dumbcow):

\[\frac{t^{2}+6t}{t-1} -\frac{7}{t-1}=\frac{t^{2}+6t-7}{t-1} = \frac{(t+7)(t-1)}{t-1} = t+7\]

OpenStudy (anonymous):

Thanks! I didn't factor the t^2 + 6t - 7, that's where I went wrong

OpenStudy (dumbcow):

no problem...always try to factor as much as you can for these type of problems